In a case of 24 plates, 3 of the plates have defects. The first plate taken from the case is defective and not returned to the case. What is the probability that the next plate sold is not defective?

1
Expert's answer
2022-03-15T03:27:23-0400

The remaining number of defective plates will be (3-1) =2 out of the remaining 23 plates.

The probability that the next plate sold is defective = 2/23

The probability that the next plate sold is not defective = (1 - 2/23)= 21/23